OpenStudy (anonymous):

Pleaseee help?? Using numerical values, prove the identity. x^3+y^3=(x+y)(x^2-xy+y^2)

OpenStudy (bibby):

assign x and y numbers and evaluate

OpenStudy (anonymous):

How do I do that?

OpenStudy (bibby):

x=2, y=5 etc/

OpenStudy (anonymous):

What are the numbers? Do i pick them myself?

OpenStudy (bibby):

I'm pretty sure you get to make them up

OpenStudy (anonymous):

Ok hold on.

OpenStudy (anonymous):

(2)^3+(3)^3=((2)+(3))((2)^2-(2)(3)+(3)^2)

OpenStudy (anonymous):

x=2 and y=3

OpenStudy (bibby):

x^3+y^3=(x+y)(x^2-xy+y^2) x=2//y=3 \[\large (2)^3+(3)^3=(2+3)((2)^2-(2)(3)+(3)^2)\] evaluate the left side now

OpenStudy (anonymous):

16+81

OpenStudy (anonymous):

@bibby you there?

OpenStudy (bibby):

420

OpenStudy (bibby):

choochin atm

OpenStudy (anonymous):

I really need help though :(

OpenStudy (bibby):

yeah let me see

OpenStudy (bibby):

(x+y)(x^2-xy+y^2) (5)(4-6+9)=(5)(7)=35

OpenStudy (bibby):

whoops

OpenStudy (anonymous):

Lol i did the left side which was 16+81

OpenStudy (bibby):

yeah, it doesn't seem to match up. I'll try it with 2 different numbers because wolfram agrees it's true

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I mean there is two sides to the problem, so dont you need a letter to bring it over?

OpenStudy (bibby):

x^3+y^3=(x+y)(x^2-xy+y^2) x=3, y=4 3^3+4^3=27+64=91 (3+4)(3^2-(3*4)+4^2) 7(9-12+16)=7*13=91

OpenStudy (anonymous):

Soooooo, is that the answer?

OpenStudy (bibby):

there is no answer. you have to pick out any 2 numbers and prove that both sides evaluate to the same number as I just did

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhh!!!!!!

OpenStudy (anonymous):

Can i ask you an other question pkease??

OpenStudy (bibby):

I'm married, sorry

OpenStudy (anonymous):

oh golly. I mean like math haha

OpenStudy (bibby):

rofl I know, post another question and tag me in it. Or post it here, I'm not a cop. the general etiquette is one question per post but i don't care about medals

OpenStudy (anonymous):

Ohhhh sowwy!

OpenStudy (bibby):

y sowwy just ask yo question

OpenStudy (anonymous):

create a polynomial of degree 3 or higher that has three or more terms. @bibby

OpenStudy (bibby):

first learn me miss professor lalalala0726 what is the degree of a polynomial

OpenStudy (anonymous):

i have no clue:( and reallyneed to get this done in like 10 minutes.. I promise ill learn!

OpenStudy (bibby):

timed test?

OpenStudy (bibby):

whatever, the degree of a polynomial is the largest exponent. so x^4 + x^3 is of degree 4

OpenStudy (bibby):

so write one out degree 3 or greater with 3 terms or more

OpenStudy (anonymous):

Ohhhhok soooo, can i do x^3+x^5?

OpenStudy (bibby):

that works. in general we sort downwards so x^5 would cum first then there's the fact that it has to be 3 terms or more and that's 2

OpenStudy (anonymous):

ok, sooo can i do x^5+x^2?

OpenStudy (bibby):

that's still 2 terms

OpenStudy (anonymous):

then instead of 2 can i do 3?

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

you kinda have to do 3 or more, add a term

OpenStudy (anonymous):

ok so x^5+x^4

OpenStudy (bibby):

that's still 2 terms

OpenStudy (anonymous):

x^5+x^4+x^3

OpenStudy (bibby):

nice

OpenStudy (bibby):

correct*